OpenStudy (anonymous):

Find the solution to the equation log16 32 = x + 3. Using complete sentences, explain the procedure used to solve this equation.

OpenStudy (anonymous):

write in exponential form: 16^(x+3) = 32 now write both sides of the equal sign as exponentials with base 2.... (2^2)^(x+3) = 2^5 2^(2x+6) = 2^5 since the bases are the same, the exponents must be the same: 2x+6 = 5 can you solve from here?

OpenStudy (anonymous):

I thin k you made a mistake here should be (2^4)^(x+3)=2^5 giving 4x+12 = 5 and x= -7/4 as 16^((-7/4)+3) = 32